OpenStudy (anonymous):

What is the solution to the compound inequality? -8<=2x-4<2 a.-6<=x<4 b. -2<=x<-1 c. 0<=x<5 d. -2<=x<3

OpenStudy (kendricklamar2014):

\[−8≤2x−4<2\] \[−8+4≤2x−4+4<2+4 ~(Add~4)\] \[−4≤2x<6\]

OpenStudy (kendricklamar2014):

\[\left(\begin{matrix}-4 \\ 2\end{matrix}\right) \le \left(\begin{matrix}2x \\ 2\end{matrix}\right)<\left(\begin{matrix}6 \\ 2\end{matrix}\right)\]

OpenStudy (kendricklamar2014):

\[Answer: −2≤x<3\]

OpenStudy (anonymous):

thank you

OpenStudy (kendricklamar2014):

You're Welcome